suppose you bought some .32 cent stamps and some .20 cent stamps. You spent $3.92 for sixteen stamps. How many of each stamp did you buy?
Answer by rwm(914)

You can't buy .32 cent stamps nor .20 cent stamps.
You can buy $.32 stamps or 32 cent stamps though.
x+y=16
32x+20y=392
x=6 (32 cent)
y=10 (20 cent)
check
6*32+10*20=392
192+200=392
ok